I have five more brothers than sisters, and each of my brothers also has five more brothers than sisters. If my family has 10 children, how many girls are there?
Thanks for the help

Found 2 solutions by solver91311, ankor@dixie-net.com:
Answer by solver91311(24713)



8 boys, 2 girls. Whichever "I" is must be one of the boys. Any one of the boys has 7 brothers which is 5 more than 2.

John

My calculator said it, I believe it, that settles it

Answer by ankor@dixie-net.com(22740)

I have five more brothers than sisters, and each of my brothers also has five more brothers than sisters.
If my family has 10 children, how many girls are there?
:
Since the speaker has the same relationship to the sisters as his brothers he is a male
Since the total is 10, the speaker has total of 9 sisters and brothers
:
"I have five more brothers than sisters,"
b = s + 5
and he has a total of 9 brothers and sisters
s + b = 9
replace b with (s+5)
s + (s+5) = 9
2s = 9 - 5
2s = 4
s = 2 sisters, two girls in the family